# Structured Sparse Transition Matrices to Enable State Tracking in State-Space Models

### v2

@@ Abstract
Modern state-space models (SSMs) often utilize transition matrices which enable efficient computation but pose restrictions on the model’s expressivity, as measured in terms of the ability to emulate finite-state automata (FSA). While unstructured transition matrices are optimal in terms of expressivity, they come at a prohibitively high compute and memory cost even for moderate state sizes. We propose a structured sparse parametrization of transition matrices in SSMs that enables FSA state tracking with optimal state size and depth, while keeping the computational cost of the recurrence comparable to that of diagonal SSMs. Our method, PD-SSM, parametrizes the transition matrix as the product of a column one-hot matrix ($P$) and a complex-valued diagonal matrix ($D$). Consequently, the computational cost of parallel scans scales linearly with the state size. Theoretically, the model is BIBO-stable and can emulate any $N$-state FSA with one layer of dimension $N$ and a linear readout of size $N\times N$, significantly improving on all current structured SSM guarantees. Experimentally, the model significantly outperforms a wide collection of modern SSM variants on various FSA state tracking tasks. On multiclass time-series classification, the performance is comparable to that of neural controlled differential equations, a paradigm explicitly built for time-series analysis. Finally, we integrate PD-SSM into a hybrid Transformer-SSM architecture and demonstrate that the model can effectively track the states of a complex FSA in which transitions are encoded as a set of variable-length English sentences. The code is available at https://github.com/IBM/expressive-sparse-state-space-model
@@ 1 Introduction
The Transformer [Vaswani et al. 2017] marked a paradigm shift in machine learning, providing a unified yet versatile architecture with strong scalability. However, experimental and theoretical results indicate that Transformers struggle in algorithmic state tracking tasks such as emulating finite-state automata (FSAs) [Hahn 2020, Bhattamishra et al. 2020, Merrill and Sabharwal 2023, Delétang et al. 2023, Liu et al. 2023, Strobl et al. 2024]. Moreover, the inherent quadratic cost of the attention mechanism hinders the training on very long sequences, and its linearly growing key/value cache can lead to memory issues during inference.
@@
In contrast, state-space models (SSMs) [Gu et al. 2022b, Gupta et al. 2022, Fu et al. 2023, Smith et al. 2023, Orvieto et al. 2023, Gu and Dao 2023] offer a scalable alternative to Transformers with complementary properties: they generalize better to longer sequences, the number of total operations scales linearly with the sequence length, and memory is constant during inference. Modern large language models (LLMs) are increasingly adopting a hybrid architecture that integrates both Transformer and SSM layers [Yang et al. 2024, Ren et al. 2025, De et al. 2024, Waleffe et al. 2024, Lenz et al. 2025, Wu et al. 2025]. To achieve computational efficiency competitive with that of Transformers, modern SSMs often utilize diagonal transition matrices. Empirically, in time-invariant SSMs, this strategy has proven effective in a range of applications [Gupta et al. 2022, Gu et al. 2022a, Orvieto et al. 2023], while time-varying diagonal SSMs can achieve language modeling performance comparable to that of the Transformer [Gu and Dao 2023].
@@
Despite their promising performance in language modeling, diagonal SSMs are restricted in the type of finite-state automaton (FSA) they can model [Merrill et al. 2024]. This stands in stark contrast with the nonlinear RNN, which tends to quickly learn such algorithmic abstractions [Delétang et al. 2023]. To enable FSA state tracking, previous approaches propose using transition matrices of various structures such as fully unstructured (i.e., dense)  [Terzić et al. 2025], or (semi-)structured matrices such as block-diagonal [Fan et al. 2024], diagonal plus low-rank (DPLR) [Schlag et al. 2021a, Yang et al. 2024, Walker et al. 2025, Peng et al. 2025] or products of DPLR matrices Siems et al. [2025]. While any $N$-state FSA can be represented by an unstructured SSM with a single layer, state dimensionality $N$, and a readout size $N\times N$, unstructured SSMs scale unfavorably and thus prohibit large-scale training. Current semi-structured approaches are more efficient, but do not allow for such compact encodings. With this in mind, the goal of this work can be summarized as follows:
@@
We aim to enable single-layer time-variant SSMs with state size $N$ and readout size $N\times N$ to represent any FSA with $N$ states while keeping the cost comparable to that of diagonal SSMs.
@@
To this end, we propose a novel structured sparse matrix parametrization of SSM transition matrices, denoted as PD parametrization, defined as a product of a binary column-one-hot matrix ($P$) and a complex-valued diagonal matrix ($D$). A PD parametrization of transition matrices is preserved under multiplication, meaning that the product of PD matrices is still a PD matrix. As such, a chained product of $L$ PD matrices can be computed using parallel scans [Blelloch 1990, Martin and Cundy 2018, Smith et al. 2023] in $\Theta(LN)$ concurrent operations and $\Theta(LN)$ memory, matching the complexity and concurrency of a chain of diagonal matrix products. In terms of expressivity, the PD matrices encompass the commonly used diagonal transition matrices, but can also represent transition functions of arbitrary $N$-state FSAs. Our contributions are as follows:
@@
  • We propose PD-SSM, a time-variant SSM that can emulate arbitrary $N$-state FSAs with a single layer, state dimension $N$, and a linear readout of size $N\times N$. Theoretically, we prove that PD-SSM is stable and that it achieves universal FSA emulation with (almost) the least possible state size.
@@
  • Empirically, we demonstrate that PD-SSM learns state tracking of complex FSAs by exhibiting state-of-the-art length generalization. We furthermore test our model on long-range multivariate time-series classification tasks and observe that it exhibits high performance, comparable to that of neural controlled differential equations.
  • We provide a novel benchmark in which transitions corresponding to FSAs are obscured by the semantics of natural language. Adopting a hybrid Transformer-SSM architecture, we empirically show that integrating PD-SSM solves the task and generalizes to longer sequences, whereas integration of diagonal SSMs fails to solve the task, even on in-domain sequences.
@@ 2.1 State-Space Models
with $u_{t}\in\mathbb{R}^{D}$ the input to the system, $x_{t}\in\mathbb{C}^{N}$ its (hidden) state, $y_{t}\in\mathbb{C}^{D}$ the complex-valued output, and $o_{t}\in\mathbb{R}^{D}$ the real-valued output. $A(u_{t})\in\mathbb{C}^{N\times N}$ is the state transition matrix, $B(u_{t})\in\mathbb{C}^{D\times N}$ and $C(u_{t})\in\mathbb{C}^{D\times N}$ map from embeddings to states and vice versa, and $D(u_{t})\in\mathbb{C}^{D\times D}$ is a memoryless skip connection. Finally, $\psi:\mathbb{C}^{D}\to\mathbb{R}^{D}$ extracts a real-valued embedding from a complex-valued embedding, often implemented as $\psi(y_{t})=Re\{y_{t}\}$ [Gu et al. 2022a, Gupta et al. 2022, Orvieto et al. 2023]. The time-variance of the system stems from the time dependence of $A(u_{t}),B(u_{t}),C(u_{t}),D(u_{t})$. In this work we focus on the $A(u_{t})$ matrix, it being the crucial differentiator between time-varying and time-invariant SSMs [Gu and Dao 2023, Merrill et al. 2024].
If the system is time-invariant, i.e., $A(u_{t})=A$, then its dynamics can be equivalently represented by a diagonal transition matrix ùp to an arbitrarily small perturbation of the entries of $A$ [Orvieto et al. 2023, Axler 2024] Indeed, writing $A=W\Lambda W^{-1}$ results, upon change of basis $\tilde{x}_{t}:=W^{-1}x_{t}$, in an equivalent state evolution $\tilde{x}_{t}=\Lambda\tilde{x}_{t}+W^{-1}Bu_{t}$ but with a diagonal state transition matrix $\Lambda$.
In contrast, if $A(u_{t})$ is a function of time, then the system admits a diagonal representation if and only if all $A(u_{t})$ are diagonalizable under the same change of basis. That is, there must exist a single $W\in\mathbb{C}^{N\times N}$ such that for all $t$, $A(u_{t})=W\Lambda_{t}W^{-1}$. This is a much more restrictive condition. It can in fact only be fulfilled if the matrix product commutes for all state transition matrices occurring over time, i.e., $\forall i\neq j,A(u_{i})A(u_{j})=A(u_{j})A(u_{i})$ [Axler 2024]. This fact hints at the restrictiveness of diagonal transition matrices, which is expanded upon in an alternative framework in Section 2.3
@@
Table 1: The structure of the transition matrix $A(u_{t})$ in time-varying SSMs determines the class of FSA they can emulate. We consider two classes of FSAs, partitioned based on the solvability of their transformation group. indicates that the structure enables the emulation of any $N$-state FSA from the partition using one layer, state dimension $N$ and a linear readout of dimension $N\times N$. * indicates that the matrix structure can emulate any automaton from the partition, but at potentially high dimension or depth, and for DPLR exponentially large linear layers. The indicated cost only considers the complexity of the recurrence. DPLR methods trade off FLOPs for increased parallelism and more optimal memory management. In the recurrent case the complexity is $\Theta(LN^{2})$, but modern methods utilize parallel algorithms which scale quadratically in $L$ [Yang et al. 2024].
@@ 2.2 Modeling Finite-State Automata with SSMs
In this work, we measure the expressivity of an SSM by considering the type of deterministic finite-state automaton (FSA) it can emulate. A deterministic FSA is an abstract model of computation, typically defined as a 5-tuple $(Q,\Sigma,\delta,q_{\text{init}},F)$ where $Q$ is a finite set of states, $\Sigma$ is a finite input alphabet, $\delta:Q\times\Sigma\rightarrow Q$ is the state transition function, $q_{\text{init}}\in Q$ is a fixed initial state, and $F\subseteq Q$ is the set of accepting states.
@@
Any (deterministic) FSA can be mapped to a time-variant SSM as follows. Encode each $q\in Q$ using $enc:Q\rightarrow\mathbb{R}^{|Q|}$ such that the encodings of different states are orthogonal. Note that orthogonality is a sufficient but not necessary condition for mapping an FSA to a selective SSM, as certain FSAs such as modular counters allow alternative, more compact, mappings to SSMs. Given such an orthogonal encoding of states, we can map the state transition function $\delta:Q\times\Sigma\rightarrow Q$ to the state transition matrices $A(u_{t})$ via $A(\sigma)={\scriptstyle\sum\nolimits}_{q\in Q}\ enc(\delta(q,\sigma))\cdot enc(q)^{T}$, set $B=0$, $C=I$, $D=0$, and $\psi=id$. Upon identification of $q$ with $enc(q)$ and $q_{\text{init}}$ with $x_{0}$, the SSM corresponds exactly to the FSA.
2.3 Limitations of Time-Variant SSMs for State Tracking in FSAs
@@
To demonstrate the limitations of diagonal SSMs for state tracking in FSAs, it suffices to consider FSAs with group structure in their state transitions. The transformation group [Straubing 1994] of such an automaton is then the algebraic group of invertible state-to-state mappings with function composition as a binary operation [^1] A more complete background on algebra and results on SSM expressivity is provided in Appendix A.. For conciseness, we equate such automata with their transformation group, thus saying, for instance, solvable automaton.
@@
Through a circuit complexity argument, Merrill et al. [2024] showed that non-solvable automata cannot be emulated using bounded-depth logarithmic precision diagonal time-variant SSMs, providing an upper bound on their expressivity. Through an explicit construction Sarrof et al. [2024] demonstrate that all solvable automata can be emulated using finite-precision complex-valued diagonal time-variant SSMs, although the depth of the SSM stack exhibits a non-trivial dependence on the complexity of the automaton. Concretely, the SSM stack depth is proportional to the Krohn-Rhodes complexity of the transformation group Margolis et al. [2024]. A more recent family of time-variant SSM models utilizes (products of) DPLR transition matrices Grazzi et al. [2024], Peng et al. [2025], Siems et al. [2025]. While such matrices enable the emulation of any FSA, the necessary SSM depth is again proportional to the Krohn-Rhodes complexity of the automaton’s transformation group Siems et al. [2025]. In contrast, unstructured SSMs Terzić et al. [2025], Walker et al. [2025] enable the emulation of any $N$-state FSA using a single layer, a state size of $N$ and a readout size $N\times N$. However, they scale poorly with the state size. An overview of all of the statements is provided in Table 1.
@@ 2.4 Associative Scan for Fully Parallelizable Recurrence in SSMs
We use parallel scans to concurrently compute the states of an SSM [Blelloch 1990, Martin and Cundy 2018, Smith et al. 2023]. Suppose a binary associative operator $\bullet$, i.e., $(a\bullet b)\bullet c=a\bullet(b\bullet c)$, and a sequence of $L$ elements $[a_{1},a_{2},\ldots,a_{L}]$. The scan operation (sometimes referred to as all-prefix-sum) then returns the sequence
$a_{1},(a_{1}\bullet a_{2}),\ldots,(a_{1}\bullet a_{2}\bullet\cdots\bullet a_{L})].$
Note that composition of the linear recurrence of a (time-variant) SSM $x_{t}=A_{t}x_{t-1}+b_{t}$ takes precisely the form of a parallel scan $(A_{t+1},b_{t+1})\bullet(A_{t},b_{t})\mapsto(A_{t+1}A_{t},A_{t+1}b_{t}+b_{t+1})$ with prefix elements $a_{t}=(A_{t},b_{t})$. According to Blelloch (1990, Section 1.4), associative scans can be computed in $\Theta(T_{\odot}\,L)$ total operations (i.e., in work-efficient manner) with $\Theta(\log L)$ sequential steps, where $T_{\odot}$ represents the cost of matrix-matrix multiplication. For dense matrices $A_{t}\in\mathbb{R}^{N\times N}$, $T_{\odot}=O(N^{3})$ and thus the associative scan quickly becomes prohibitively expensive in deep learning settings. In contrast, if $A_{t}$ are diagonal matrices, $T_{\odot}=O(N)$. In terms of memory, parallel scans with dense matrices use $\Theta(N^{2}L)$ memory, whereas diagonal matrices lead to a $\Theta(NL)$ memory consumption.

### v3

@@ Abstract
Modern state-space models (SSMs) often utilize structured transition matrices which enable efficient computation but pose restrictions on the model’s expressivity, as measured in terms of the ability to emulate finite-state automata (FSA). While unstructured transition matrices are optimal in terms of expressivity, they come at a prohibitively high compute and memory cost, even for moderate state sizes. We propose a structured sparse parametrization of transition matrices in SSMs that enables FSA state tracking with provably optimal state size and depth, while keeping the computational cost of the recurrence comparable to that of diagonal SSMs. Our method, PD-SSM, parametrizes the transition matrix as the product of a column one-hot matrix ($P$) and a complex-valued diagonal matrix ($D$). As a result, the computational cost of parallel scans scales linearly with the state size. Theoretically, the model is BIBO-stable and can emulate any $N$-state FSA with one layer of dimension $N$ and a linear readout of size $N\times N$, significantly improving on all current structured SSM guarantees. Experimentally, the model significantly outperforms a wide collection of modern SSM variants on various FSA state tracking tasks. On multivariate time-series classification, it outperforms neural controlled differential equations, a paradigm explicitly built for time-series analysis. Finally, we integrate PD-SSM into a hybrid Transformer-SSM architecture and demonstrate that the model can effectively track the states of a complex FSA in which transitions are encoded into sets of variable-length English sentences. The code is available at https://github.com/IBM/expressive-sparse-state-space-model.
@@ 1 Introduction
The Transformer [vaswani_attention_2017] marked a paradigm shift in machine learning, providing a unified yet versatile architecture with strong scalability. However, experimental and theoretical results indicate that Transformers struggle in algorithmic state tracking tasks such as finite-state automaton (FSA) emulation [hahn_2020_theoretical, bhattamishra_ability_2020, merrill_parallelism_2023, deletang_neural_2023, liu_transformers_2023, strobl_2024_survey]. Moreover, the inherent quadratic cost of the attention mechanism hinders the training on very long sequences, and its linearly growing key/value cache can lead to memory issues during inference.
@@
In contrast, state-space models (SSMs) [gu_efficiently_2022, gupta_diagonal_2022, fu_hungry_2023, smith_simplified_2023, orvieto_resurrecting_2023, gu_mamba_2023] offer a scalable alternative to Transformers with complementary properties: they generalize better to longer sequences, the number of total operations scales linearly with the sequence length, and memory is constant during inference. Modern large language models (LLMs) are increasingly adopting a hybrid architecture that integrates both Transformer and SSM layers [yang2024parallelizing, ren_2024_samba, de_griffin_2024, waleffe_2024_empirical, lenz_2025_jamba, wu2025transxssm]. To achieve computational efficiency competitive with that of Transformers, modern SSMs often utilize diagonal transition matrices. Empirically, in time-invariant SSMs, this strategy has proven effective in a range of applications [gupta_diagonal_2022, gu_parameterization_2022, orvieto_resurrecting_2023], while time-varying diagonal SSMs can achieve language modeling performance comparable to that of the Transformer [gu_mamba_2023].
@@
Despite their promising performance in language modeling, diagonal SSMs are restricted in the type of FSA they can emulate [merrill_illusion_2024]. This stands in stark contrast with the nonlinear RNN, which can quickly learn such algorithmic abstractions [deletang_neural_2023]. To enable FSA state tracking, previous approaches propose using transition matrices of various structures such as fully unstructured (i.e., dense)  [terzic2025sdssm], or (semi-)structured matrices such as block-diagonal [fan_advancing_2024], diagonal plus low-rank (DPLR) [schlag2021linear, yang2024parallelizing, walker_2025_structured, peng_2025_rwkv] or products of DPLR matrices siems_2025_deltaproduct. While any $N$-state FSA can be represented by an unstructured SSM with a single layer, state dimensionality $N$, and a readout size $N\times N$, unstructured SSMs scale unfavorably and thus prohibit large-scale training. Current (semi-)structured approaches are more efficient, but do not allow for such compact encodings. With this in mind, the goal of this work can be summarized as follows:
@@
We aim to enable single-layer time-varying SSMs with state size $N$ and readout size $N\times N$ to represent any FSA with $N$ states while keeping the cost comparable to that of diagonal SSMs.
@@
To this end, we propose a novel structured sparse matrix parametrization of SSM transition matrices, denoted as PD parametrization, defined as a product of a binary column-one-hot matrix ($P$) and a complex-valued diagonal matrix ($D$). A PD parametrization of transition matrices is preserved under multiplication, meaning that the product of PD matrices is still a PD matrix. As such, a chained product of $L$ PD matrices can be computed using parallel scans [blelloch_prex_1990, martin_parallelizing_2018, smith_simplified_2023] in $\Theta(LN)$ concurrent operations and $\Theta(LN)$ memory, matching the complexity and concurrency of a chain of diagonal matrix products. In terms of expressivity, the PD matrices encompass the commonly used diagonal transition matrices, but can also represent transition functions of arbitrary $N$-state automata. Our contributions are as follows:
@@
  • We propose PD-SSM, a time-varying SSM that can emulate any $N$-state FSA with a single layer, state dimension $N$, and a linear readout of size $N\times N$. Theoretically, we prove that PD-SSM is stable and that it achieves universal FSA emulation with (almost) the least possible state size.
@@
  • Empirically, we demonstrate that PD-SSM learns to track the states of complex automata by exhibiting state-of-the-art length generalization. We furthermore test our model on long-range multivariate time-series classification tasks and observe that it reaches state-of-the-art performance, even surpassing neural controlled differential equations.
  • We provide a novel benchmark in which each FSA transition is redundantly encoded by meaningful English sentences. Adopting a hybrid Transformer-SSM architecture, we show that integrating PD-SSM enables the model to emulate the underlying FSA and to generalize to longer sequences, in contrast to diagonal SSMs which fail on in-domain sequences.
@@ 2.1 State-Space Models
with $u_{t}\in\mathbb{R}^{D}$ the input to the system, $x_{t}\in\mathbb{C}^{N}$ its (hidden) state, $y_{t}\in\mathbb{C}^{D}$ the complex-valued output, and $o_{t}\in\mathbb{R}^{D}$ the real-valued output. $A(u_{t})\in\mathbb{C}^{N\times N}$ is the state transition matrix, $B(u_{t})\in\mathbb{C}^{D\times N}$ and $C(u_{t})\in\mathbb{C}^{D\times N}$ map from embeddings to states and vice versa, and $D(u_{t})\in\mathbb{C}^{D\times D}$ is a memoryless skip connection. Finally, $\psi:\mathbb{C}^{D}\to\mathbb{R}^{D}$ extracts a real-valued embedding from a complex-valued embedding, often implemented as $\psi(y_{t})=Re\{y_{t}\}$ [gu_parameterization_2022, gupta_diagonal_2022, orvieto_resurrecting_2023]. The time-variance of the system stems from the time dependence of $A(u_{t}),B(u_{t}),C(u_{t}),D(u_{t})$. In this work we only consider time dependence in the $A(u_{t})$ matrix, this being the crucial differentiator between time-varying and time-invariant SSMs [gu_mamba_2023, merrill_illusion_2024]. We thus fix $B(u_{t})=B,C(u_{t})=C,D(u_{t})=D$.
[…]
In contrast, if $A(u_{t})$ is a function of time, then the system admits a diagonal representation if and only if all $A(u_{t})$ are diagonalizable under the same change of basis. That is, there must exist a single $W\in\mathbb{C}^{N\times N}$ such that for all $t$, $A(u_{t})=W\Lambda_{t}W^{-1}$. This is a much more restrictive condition. It can in fact only be fulfilled if the matrix product commutes for all state transition matrices occurring over time, i.e., $\forall i\neq j,A(u_{i})A(u_{j})=A(u_{j})A(u_{i})$ [axler_linear_2024]. This fact hints at the restrictiveness of diagonal transition matrices, which is expanded upon in an alternative framework in Section 2.3
@@
Table 1: The structure of the transition matrix $A(u_{t})$ in time-varying SSMs determines the class of FSA they can emulate. We consider two classes of automata, partitioned based on the solvability of their transformation group. indicates that the structure enables the emulation of any $N$-state FSA from the partition using one layer, state dimension $N$ and a linear readout of dimension $N\times N$. * indicates that the matrix structure can emulate any automaton from the partition, but at potentially high dimension or depth, and in case of RWKV-7 exponentially large linear layers siems_2025_deltaproduct. The cost only considers the complexity of the parallel computation of $x_{t}$. Modern DPLR methods trade off FLOPs for increased parallelism and better memory management. The recurrent complexity is $\Theta(LN^{2})$, but modern methods utilize algorithms that scale quadratically in $L$ [yang2024parallelizing].
@@ 2.2 Modeling Finite-State Automata with SSMs
In this work, we measure the expressivity of an SSM by considering the type of deterministic finite-state automaton (FSA) it can emulate. A deterministic FSA is an abstract model of computation defined as a 5-tuple $(Q,\Sigma,\delta,q_{\text{init}},F)$ where $Q$ is a finite set of states, $\Sigma$ is a finite input alphabet, $\delta:Q\times\Sigma\rightarrow Q$ is the state transition function, $q_{\text{init}}\in Q$ is a fixed initial state, and $F\subseteq Q$ is the set of accepting states.
@@
Any (deterministic) FSA can be mapped to a time-variant SSM as follows. Encode each $q\in Q$ using $enc:Q\rightarrow\mathbb{R}^{|Q|}$ such that the encodings of different states are orthogonal. Note that orthogonality is a sufficient but not necessary condition for mapping an FSA to a selective SSM, as certain automata such as modular counters allow alternative, more compact, mappings to SSMs. Given such an orthogonal encoding of states, we can map the state transition function $\delta:Q\times\Sigma\rightarrow Q$ to the state transition matrices $A(u_{t})$ via $A(\sigma)={\scriptstyle\sum\nolimits}_{q\in Q}\ enc(\delta(q,\sigma))\cdot enc(q)^{T}$, set $B=0$, $C=\mathbb{I}$, $D=0$, and $\psi=id$. Upon identification of $q$ with $enc(q)$ and $q_{\text{init}}$ with $x_{0}$, the SSM matches the FSA exactly.
2.3 Limitations of Time-Variant SSMs for FSA State Tracking
@@
To demonstrate the limitations of various SSMs structures for state tracking in automata, it suffices to consider automata with group structure in their state transitions. The transformation group [straubing_book] of such an automaton is then the algebraic group of invertible state-to-state mappings with function composition as a binary operation [^2] A more complete background on algebra and results on SSM expressivity is provided in Appendix A.. For conciseness, we equate such automata with their transformation group, thus saying, for instance, solvable automaton.
@@
Through a circuit complexity argument, merrill_illusion_2024 showed that non-solvable automata cannot be emulated using bounded-depth logarithmic precision diagonal time-variant SSMs, providing an upper bound on their expressivity. Through an explicit construction sarrof2024expressivecapacitystatespace demonstrate that all solvable automata can be emulated using finite-precision complex-valued diagonal time-variant SSMs, although the depth of the SSM stack exhibits a non-trivial dependence on the complexity of the FSA. Concretely, the SSM stack depth is proportional to the Krohn-Rhodes complexity of the transformation group margolis_2024_decidability. A more recent family of time-variant SSM models utilizes diagonal plus low-rank (DPLR) transition matrices grazzi2024unlocking, peng_2025_rwkv, siems_2025_deltaproduct. While such matrices enable the emulation of any FSA, the necessary model depth is either a function of the automaton’s structure, or the model requires exponentially large linear layers siems_2025_deltaproduct. In contrast, unstructured SSMs terzic2025sdssm, walker_2025_structured enable the emulation of any $N$-state FSA using a single layer, a state size of $N$ and a readout size $N\times N$. However, they scale poorly with the state size. For an overview of the results, see Table 1.
@@ 2.4 Associative Scan for Fully Parallelizable Recurrence in SSMs
We use parallel scans to concurrently compute the states of an SSM [blelloch_prex_1990, martin_parallelizing_2018, smith_simplified_2023]. Suppose a binary associative operator $\bullet$, i.e., $(a\bullet b)\bullet c=a\bullet(b\bullet c)$, and a sequence of $L$ elements $[a_{1},a_{2},\ldots,a_{L}]$. The scan operation (sometimes referred to as all-prefix-sum) then returns the sequence $[a_{1},(a_{1}\bullet a_{2}),\ldots,(a_{1}\bullet a_{2}\bullet\cdots\bullet a_{L})].$ Note that composition of the linear recurrence of a (time-variant) SSM $x_{t}=A_{t}x_{t-1}+b_{t}$ takes precisely the form of a parallel scan $(A_{t+1},b_{t+1})\bullet(A_{t},b_{t})\mapsto(A_{t+1}A_{t},A_{t+1}b_{t}+b_{t+1})$ with prefix elements $a_{t}=(A_{t},b_{t})$. According to blelloch_prex_1990 (blelloch_prex_1990, Section 1.4), associative scans can be computed in $\Theta(T_{\odot}\,L)$ total operations (i.e., in work-efficient manner) with $\Theta(\log L)$ sequential steps, where $T_{\odot}$ represents the cost of matrix-matrix multiplication. For dense matrices $A_{t}\in\mathbb{R}^{N\times N}$, $T_{\odot}=O(N^{3})$ and thus the associative scan quickly becomes prohibitively expensive in deep learning settings. In contrast, if $A_{t}$ are diagonal matrices, $T_{\odot}=O(N)$. In terms of memory, parallel scans with dense matrices use $\Theta(N^{2}L)$ memory, whereas diagonal matrices lead to a $\Theta(NL)$ memory consumption.